OpenStudy (anonymous):

Simplify answer: 25/3 x 7/5=

OpenStudy (anonymous):

35/3

OpenStudy (anonymous):

\[\frac{25}{3}*\frac{7}{5}=\frac{5*5*7}{5*3}=\frac{5*7}{3}=\frac{35}{3}\]

OpenStudy (anonymous):

wow thank you

OpenStudy (anonymous):

is it clear how that works ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Click "Good Answer" for helpful answers :)